if an economy is growing at 10% , how long it will take for $1000 turn into $4000?
Answer #1

Answer

By the rule of 70

year to double the economy =70/growth rate

=70/10

=7 years

1000 to 4000 is 4 times

so

it doubles in 7 years and double of double in 14 years so the economy will take 14 years for 1000 to tern 4000.

======

By formula

n=ln(FV/PV)/(ln(1+i)

FV=future value =4000

PV=present value =1000

i=10%

n=years

n=ln(4000/1000)/ln(1+0.1)

=14.5450818

=14.5

it will take 14.5 years